Circular No. 7975

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

ANOTHER NOVA IN SAGITTARIUS Katsumi Haseda, Aichi, Japan, reports his discovery of a possible nova

(mag ∼ 5.0) on Tri-X film taken with a 120-mm f/3.5 camera lens on Sept. 20.431 UT; Haseda provided an approximate position for the new star of α = 19^{h}01^{m}08^{s}.5, δ = −22^{o}00′45′′ (equinox 2000.0). D. West, Mulvane, KS, reports the following precise position end figures measured from a CCD exposure taken on Sept. 21.07: 09^{s}.38, 05′′.9 (uncertainty ± 0′′.75; twenty GSC 1.1 reference stars). West adds that nothing is visible at this location on a Digitized Sky Survey image from 1987.59. H. Yamaoka, Kyushu University, and G. Masi, Ceccano, Italy, report that a USNO-A2.0 star is located at position end figures 09^{s}.321, 06′′.17 with red mag 16.7 and blue mag 17.0. Yamaoka adds that the second-nearest object in the same catalogue is ∼ 7′′ from West’s position.

T. Kato, Kyoto University, forwards a report from M. Fujii (Fujii Bisei Observatory, Okayama, Japan; (via K. Ayani, Bisei Astronomical Observatory) that a spectrum of this new object, taken on Sept. 21.44, shows strong Balmer and Fe II emission lines and confirms that the object is an Fe II-class nova shortly after maximum; the FWHM of Hα emission is 2400 km/s.

Magnitude estimates, visual unless otherwise noted: Sept. 19.424, 5.5: (T. Tanaka, Mie, Japan, FinePix S2Pro unfiltered CCD camera; magnitude estimated by Kato); 19.427, 4.9 (H. Nishimura, Kakegawa, Shizuoka, Japan, T-Max 400 film; communicated by Kato); 21.07, B = 6.44 (West; CCD); 21.145, 6.1 (J. Garcia, Rama Caída, Argentina); 21.229, 5.8 (M. Linnolt, Honolulu, HI); 21.267, 5.8 (J. Bedient, Honolulu, HI); 21.409, 5.8 (Linnolt); 21.33, 5.5 (S. J. O’Meara, Volcano, HI).

COMET C/2002 S2 (SOHO) Further to IAUC 7969, M. Uzzo reports his measurements for a Kreutz

sungrazer found by S. Hönig on SOHO website C3 coronagraph images. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-S17.

2002 UT  α_{2000}  δ_{2000}
Sept.17.571  11^{h}06.^{m}7  + 0^{o}39′
2002 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green